OpenStudy (anonymous):

What is the solution to the set of lx-2l>3? X<1 or x>-5 X<-1 or x>5 X>1 or x<-5 X>-1 or x<5

OpenStudy (anonymous):

What is the degree of the polynomial 2h^5+3h^4+h^2-3

OpenStudy (anonymous):

1. \[\left| x-a \right|>b \implies \left( x-a \right) <-b ~and~\left( x-a \right)>b\]

OpenStudy (anonymous):

2. highest degree of h

OpenStudy (anonymous):

I have no clue what you said

OpenStudy (solomonzelman):

do you know what |x-2| means ?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

I also have 1 more question after this if you don't mind helping

OpenStudy (solomonzelman):

we can't jump like this (it's not basketball)

OpenStudy (solomonzelman):

So, when you say |4| that means "absolute value of a 4 (i.e. a distance of 4)

OpenStudy (solomonzelman):

tell, me what do you get for 4-0 and for 4-8 ?

OpenStudy (anonymous):

4 and -4

OpenStudy (solomonzelman):

yes

OpenStudy (anonymous):

Do I just plug in the numbers into x?

OpenStudy (solomonzelman):

hold on, follow along with me

OpenStudy (anonymous):

Okay

OpenStudy (solomonzelman):

if you say |4-0| or |4-8| (the first one mean "the absolute value of 4 minus 0" , and that is "the distance from 4 to 0") (the first one mean "the absolute value of 4 minus 8" , and that is "the distance from 4 to 8") both of these distances (from 4 to 0, and from 4 to 8) are (each of them is) =4

OpenStudy (solomonzelman):

let me draw this

OpenStudy (solomonzelman):



OpenStudy (solomonzelman):